OpenStudy (anonymous):

a parabola opens upward and has a vertex at (-4,3). which of the following could represent the parabola?

OpenStudy (anonymous):

the one that looks like \[y=a(x+4)^2+3\]

OpenStudy (anonymous):

i dont know it just says what i posted.

OpenStudy (anonymous):

answeres are: A. f(x)=(x-4)^2+3. B.f(x)=(x+4)^2-3. C. f(x)=(x+4)^2+3. D. f(x)=(x-4)^2-3.

OpenStudy (anonymous):

did you not get the answer i wrote?

OpenStudy (anonymous):

srry! yes i did! oops!

OpenStudy (anonymous):

whew in this case \(a=1\)

OpenStudy (anonymous):

its C dude

OpenStudy (anonymous):

thanks. i see that.